How many grams of fluorine gas are in 3.28×10 to the fifth power L at a pressure of 5.79 kpa and a temperature of 46 degree celsius?
PV=nRT
Use R=0.80206 L atm/k mol
1
Expert's answer
2020-03-23T10:15:51-0400


How many grams of fluorine gas are in 3.28×105 L at a pressure of 5.79 kpa and a temperature of 46 o C?

PV=nRT

Use R=0.080206 L atm/k mol


Solution:

According to Mendeleev-Klapeyrone's equation:

pV = nRT (1)

or:

pV = mRT/M, where m is mass in grams; M - molar mass of substance in g/mol.

M(F2) = 2*19 = 38 g/mol

Using equation (1) we can calculate the mass of F2:

m(F2) = pV*M/(RT)

where p is pressure which equals 5.79 kpa = 0.0571 atm;

T is temperature = 46oC = 319 K

Thus:

m(F2) = 0.0571*3.28*105*38/(0.080206*319) = 27816.1 (g) = 27.816 (kg)


Answer: m(F2) = 27816.1 grams.